# Developmentally-plausible Working Memory Shapes a Critical Period for Language Acquisition

### v1

@@ 4.2 Baselines
  • NoLimit: A model with no memory constraints. Working memory remains constant from the early stages of training, simulating the mature working memory observed post-adolescence. This configuration is equivalent to a vanilla GPT-2 model.
@@
Model Overall D-n agr S-v agr Ana. agr Arg. str Binding Case Ellipsis Filler. gap Irregular Island Local. atr Quantifiers NPI AO-CHILDES DynamicLimit (↑) 62.2 50.8 50.0 49.6 67.7 58.7 95.2 43.1 93.6 52.2 53.6 51.3 57.6 85.0 DynamicLimit (↓) 56.5 49.9 49.7 50.1 44.7 61.9 70.6 73.3 72.0 51.8 61.9 47.0 48.1 54.1 $\Delta$ (↑, ↓) 5.7 0.9 0.3 -0.5 23.0 -3.2 24.6 -30.1 21.6 0.4 -8.3 4.4 9.5 30.8 Wikipedia DynamicLimit (↑) 59.1 49.8 50.4 46.0 71.5 59.3 97.7 37.4 86.5 51.1 58.0 60.5 42.2 53.9 DynamicLimit (↓) 52.9 50.4 50.1 47.4 68.7 62.3 74.4 60.2 44.2 53.2 51.7 42.7 40.6 42.2 $\Delta$ (↑, ↓) 6.1 -0.6 0.3 -1.4 2.9 -3.0 23.3 -22.8 42.3 -2.2 6.3 17.8 1.7 11.7
Table 3: Performance difference when changing the direction of the cognitive constraints of DynamicLimit-Exp
@@ 4.3 Results
Developmentally-plausible working memory shapes the critical period for language acquisition
Table 1 presents the accuracy of each model trained on the AO-CHILDES dataset. Compared to NoLimit and StaticLimit, which do not account for developmental changes in working memory, DynamicLimit-Linear and DynamicLimit-Exp, which simulate its gradual growth, achieve significantly higher overall performance. Among them, DynamicLimit-Exp attains the highest overall accuracy, further supporting the effectiveness of a cognitively plausible mechanism.
The comparable performance of StaticLimit to NoLimit, despite incorporating head-specific static working memory constraints, suggests that the gradual introduction of working memory constraints throughout training is crucial, rather than their static application. These results indicate that DynamicLimit-Exp effectively replicates the critical period effect observed in human L1 acquisition.
The critical period depends on the child’s learning algorithm, not the input stimulus
Table 2 presents the accuracy of models trained on Wikipedia, showing trends similar to those observed in Table 1, where the models were trained on AO-CHILDES. Specifically, DynamicLimit-Linear and DynamicLimit-Exp outperform NoLimit and StaticLimit in overall accuracy, with DynamicLimit-Exp achieving the highest performance, further supporting the efficacy of incorporating developmental working memory constraints.
These findings suggest that the critical period effect does not depend solely on exposure to specific linguistic stimuli (e.g., CDS) but rather on the learning algorithm itself, which mirrors human cognitive development. This result aligns with existing research Feng et al. (2024), which has reported that child language input is not uniquely valuable for training language models. Furthermore, this finding suggests the potential applicability of modeling the gradual growth of working memory in the pretraining of LLMs trained on non-CDS datasets (e.g., Wikipedia).
@@
(b) DynamicLimit
Variance Entropy Mean Distance
Epoch 1 5 10 1 5 10 1-5 5-10 1-10
NoLimit [506.68, 2089.67] [1397.64, 2888.55] [616.50, 1620.18] 5.30 5.23 5.30 75.47 12.26 87.62
DynamicLimit-Exp [547.48, 2547.30] [732.21, 2535.31] [728.44, 1528.82] 5.29 5.30 5.34 59.91 37.68 97.59
Table 4: Embedded space analysis of each learning stage of NoLimit and DynamicLimit-Exp: cluster expansion, distribution diversity, distance between distributions. The variance indicates [variance in the x-axis direction, variance in the y-axis direction].
@@ 5.1 Testing the “Less-is-more” Hypothesis with Reversed Cognitive Constraints
An important question arising from the results is whether the superior performance of DynamicLimit-Exp is attributable to the “Less-is-more” hypothesis Newport (1990)—i.e., the gradual growth of working memory—or whether it results from an unintended side effect, such as an increase in information capacity, where the gradual change in working memory dynamically shifts the model’s focus across epochs, ultimately facilitating rule generalization. To test this, we introduce a cognitively implausible language model, referred to as “DynamicLimit-Exp (↓)”, which shares the same slope trajectory as our proposed DynamicLimit-Exp (↑)  [^7] This section adopts this notation for simplicity. but with its direction reversed, such that working memory capacity decreases over time.
The results in Table 3 provide evidence supporting the Less-is-more hypothesis, as DynamicLimit-Exp (↑) consistently outperformed the cognitively implausible DynamicLimit-Exp (↓). The observed performance gap, particularly in grammatical items requiring both local and non-local dependencies (e.g., CASE, ARG. STR, and FILLER-GAP), suggests that the gradual growth of working memory is crucial for grammatical learning and generalization, as it enables the early extraction of basic patterns followed by the progressive acquisition of complex rules. These findings indicate that the superior performance of DynamicLimit-Exp (↑) is primarily driven by the developmental trajectory of working memory growth rather than unintended side effects of dynamic shifts in memory focus.
@@ 5.2 Development of Feature Extraction Capabilities
By comparing features extracted in the final layers of NoLimit and DynamicLimit-Exp, we further analyze the hypothesis that gradual working memory growth enables the extraction of simple patterns in early training stages and more complex patterns in later stages. Figure 3 visualizes the clustering structure of final layer embeddings using t-SNE van der Maaten and Hinton (2008) for CASE, the grammatical item where gradual growth curves improved performance in both AO-CHILDES and Wikipedia, as identified in the previous analysis (§5.1). The embedding space visualizations reveal distinct patterns between NoLimit and DynamicLimit-Exp across training epochs. In NoLimit, the embedding clusters expand between Epoch 1 and Epoch 5 but contract significantly by Epoch 10, suggesting stagnation in representation learning. In contrast, DynamicLimit-Exp maintains structured evolution throughout training, with well-separated clusters that reflect progressive refinement.
To quantitatively analyze the differences in embedding spaces, Table 4 presents key characteristics of the embedding distributions, including variance (spread of clusters), entropy (distribution diversity), and mean Euclidean distance (inter-cluster distance). From the perspective of variance, NoLimit exhibits a temporary increase between Epoch 1 and Epoch 5, followed by a sharp reduction at Epoch 10, particularly along the y-axis. This pattern suggests an initial capacity to capture diverse features that diminishes over time. Conversely, DynamicLimit-Exp maintains substantial variance across epochs, indicating the model’s ability to progressively adapt its representation space.
Furthermore, the improvement in language acquisition efficiency can also be explained from the perspective of isotropy (or anisotropy) in the embedding space. From an anisotropy perspective, NoLimit exhibits increasing clustering of embeddings as training progresses, particularly by Epoch 10, leading to a highly compact representational space. This anisotropic structure suggests a limited capacity for diverse representation learning and stagnation in generalization. Conversely, DynamicLimit-Exp maintains a more isotropic distribution of embeddings throughout training, as evidenced by higher variance and sustained entropy. This isotropy ensures that embeddings are more evenly distributed across the representational space, facilitating consistent refinement and improved handling of linguistic diversity. These findings align with recent work on syntactic smoothing, which suggests that reducing anisotropy enhances the ability to generalize to infrequent tokens Diehl Martinez et al. (2024). Thus, the improved isotropy observed in DynamicLimit-Exp provides strong evidence for its capacity to support effective rule generalization and representation learning.
@@ 5.3 Influence of Input Stimulus Length
We further investigate the influence of sentence length on the performance of two models: NoLimit and DynamicLimit-Exp. To evaluate their adaptability to varying sentence lengths, we constructed four datasets, each containing 500,000 sentences with lengths in the ranges [5,10], [11,50], [51,100], and [101,150], respectively. These datasets were derived from Wikipedia to ensure consistency in data quantity across all sentence-length conditions.
The results in Table 5 reveal notable differences in model performance. For shorter sentences in the [5,10] range, NoLimit achieves slightly higher accuracy compared to DynamicLimit-Exp. However, in the [11,50] range, DynamicLimit-Exp significantly outperforms NoLimit, achieving 58.7 compared to 47.0. This suggests that DynamicLimit-Exp excels at handling moderately long sentences, likely due to its ability to dynamically adjust working memory. For longer sentences in the [51,100] and [101,150] ranges, DynamicLimit-Exp consistently outperforms NoLimit. The performance of NoLimit declines rapidly as sentence length increases, indicating its limitations in processing long-range dependencies. In contrast, DynamicLimit-Exp demonstrates greater robustness to increasing sentence length, likely due to its capacity to learn complex patterns over longer contexts.
These findings highlight the benefits of dynamic working memory expansion in facilitating rule generalization and contextual adaptation across diverse sentence lengths. While NoLimit exhibits competitive performance on short sentences, its stagnation on longer sentences underscores its limited ability to generalize complex patterns. Conversely, DynamicLimit-Exp’s consistent performance across varying sentence lengths supports its suitability for tasks requiring the processing of both short and long contexts. This analysis underscores the importance of tailoring model architecture to the specific requirements of NLP tasks. The results suggest that dynamic memory strategies, such as those employed in DynamicLimit-Exp, can significantly enhance performance, particularly in scenarios requiring the comprehension of lengthy and complex sentences.
@@ 6 Conclusion
This study proposed a method for integrating the developmental trajectory of human working memory into language models, inspired by the Less-is-More hypothesis. The proposed method, DynamicLimit-Exp, initially restricts working memory and gradually relaxes it exponentially during training. Experiments on AO-CHILDES and Wikipedia datasets showed that DynamicLimit-Exp improves grammatical learning efficiency compared to models with static or no memory constraints. Further analysis confirmed that gradual memory expansion enhances rule extraction and generalization. By testing a reversed version of our method, we provided strong evidence supporting the Less-is-More hypothesis, demonstrating that early cognitive limitations facilitate learning. These results suggest that the development of working memory may play a crucial role in shaping the critical period and provide new insights for improving data efficiency and understanding language acquisition.
@@ Acknowledgments
One limitation of this study is the scale of the experimental setup. The primary goal of this study is to computationally replicate the critical period in L1 acquisition, as discussed in cognitive science Lenneberg (1967); Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). Following previous studies Huebner et al. (2021); Haga et al. (2024), we designed the experiment to be as ecologically valid as possible by training an SLM using CDS. While this controlled setting allows for a more precise analysis and simulation of the Less-is-More hypothesis, it remains unclear how our findings contribute to the data efficiency of LLMs. The experimental results with Wikipedia (Table 2, 3, 5) provide a promising outlook in this direction, but further investigation with larger models and datasets is necessary to determine the effectiveness and limitations of the proposed approach.
@@
In this experiment, we investigated the replication of the critical period effect in L1 acquisition using English. However, since the critical period effect is observed across various languages Patkowski (1980); Johnson and Newport (1989), it remains to be tested whether the proposed approach is effective in multilingual environments. To our knowledge, there is currently no targeted syntactic evaluation specifically designed for CDS across different languages, such as Zorro. Zorro was developed based on BLiMP Warstadt et al. (2020), an adult-oriented targeted syntactic evaluation for English, and recent studies have proposed multilingual versions of BLiMP (e.g., JBLiMP Someya and Oseki (2023) for Japanese and CLiMP Xiang et al. (2021) for Chinese). Therefore, developing CDS-specific versions based on these multilingual BLiMPs could help address this limitation.
@@ Appendix A Details of the Training Configuration for the Base Models
See Table 6
@@
Maximum Sequence Length 32 (AO-CHILDES) /64 (Wikipedia)
@@
Optimizer AdamW, $\beta=(0.9,0.999)$ and $\epsilon=1e−08$
Tokenizer Trained on CHILDES Macwhinney (2000)
@@ Appendix B Details of Grammatical Items in Zorro
Table 7: Explanation of each grammatical category in Zorro. Examples are taken from Table 5 in the original paper Huebner et al. (2021).
See Table 7

### v2

@@ 4.2 Baselines
  • NoLimit: A model with no memory constraints. Working memory remains constant from the early stages of training, simulating the mature working memory observed post-adolescence. This configuration is equivalent to a vanilla GPT-2 Radford et al. (2019).
@@ 4.3 Results
Developmentally-plausible working memory shapes the CP for $\mathrm{L}_{1}$ acquisition
Table 1 presents the accuracy of each model trained on the AO-CHILDES. Compared to NoLimit and StaticLimit, which do not account for developmental changes in working memory, DynamicLimit-Linear and DynamicLimit-Exp, which simulate its gradual growth, achieve significantly higher overall performance. Among them, DynamicLimit-Exp attains the highest overall accuracy, supporting the effectiveness of a cognitively plausible mechanism. The comparable performance of StaticLimit to NoLimit suggests that the gradual introduction of working memory constraints throughout training is crucial, rather than their static application. These results indicate that DynamicLimit-Exp effectively replicates the CP effect observed in human $\mathrm{L}_{1}$ acquisition.
The CP depends on the child’s learning algorithm, not the input stimulus
Table 2 presents the accuracy of models trained on Wikipedia, showing trends similar to those observed in Table 1, where the models were trained on AO-CHILDES. Specifically, DynamicLimit-Linear and DynamicLimit-Exp outperform NoLimit and StaticLimit in overall accuracy, with DynamicLimit-Exp achieving the highest performance, further supporting the efficacy of incorporating developmental working memory constraints. These findings suggest that the CP effect does not depend solely on exposure to specific linguistic stimuli (e.g., CDS) but rather on the learning algorithm itself, which mirrors human cognitive development.
This result aligns with existing research Feng et al. (2024), which has reported that child language input is not uniquely valuable for training LMs. This finding suggests that our method is applicable to LLM pretraining, as they typically use non-CDS datasets such as Common Crawl and Wikipedia Touvron et al. (2023a).
@@
(b) DynamicLimit-Exp
Figure 3: Embedded space at each learning stage for NoLimit and DynamicLimit-Exp (Filler. gap)
@@ 5.1 Testing the “Less-is-more” Hypothesis with Reversed Cognitive Constraints
A key question arising from the results (§4) is whether DynamicLimit-Exp’s superior performance stems from the “Less-is-more” hypothesis Newport (1990)—i.e., the gradual growth of working memory—or from unintended side effects. In other words, does the gradual change in working memory enhance information capacity, dynamically shifting the model’s focus across epochs and ultimately aiding rule generalization? To test this, we introduce a cognitively implausible language model, referred to as “DynamicLimit-Exp (↓)”, which shares the same slope trajectory as our proposed DynamicLimit-Exp (↑)  [^6] This section adopts this notation for simplicity. but with its direction reversed, such that working memory capacity decreases over time. Specifically, DynamicLimit-Exp (↑) is set to $m_{0}=1.0,r=0.6$ (the same setting as in §4), while DynamicLimit-Exp (↓) is set to $m_{0}=0.01,r=1.668$ to achieve a nearly symmetrical curve. [^7] Since setting the initial slope $m_{0}=0.0$ prevents $w_{t}$ from being updated in Equation (2), we set it this way for computational reasons.
Table 3 provides evidence supporting the Less-is-more hypothesis, as DynamicLimit-Exp (↑) consistently outperformed the cognitively implausible DynamicLimit-Exp (↓). The observed performance gap, particularly in grammatical items requiring both local and non-local dependencies (e.g., CASE, ARG. STR, and FILLER-GAP), suggests that the gradual growth of working memory is crucial for grammatical learning and generalization, as it enables the early extraction of basic patterns followed by the progressive acquisition of complex rules. These findings indicate that the superior performance of DynamicLimit-Exp (↑) is primarily driven by the developmental trajectory of working memory growth rather than unintended side effects of dynamic shifts in memory focus.
Entropy Mean Distance
Epoch 1 5 10 1-5 5-10 1-10
NoLimit 5.36 5.17 5.19 91.30 28.50 66.28
DynamicLimit-Exp 5.40 5.30 5.39 69.25 70.63 101.92
Table 4: Embedded space analysis of NoLimit and DynamicLimit-Exp at each stage: distribution diversity and distribution distance.
[…]
Mark fixed one worn canal, and Roger
  fixed more.
*Mark fixed one canal, and Roger fixed
  more worn. Since resolving ELLIPSIS involves maintaining long-range dependencies, DynamicLimit-Exp (↑) may struggle due to its initial memory constraints. This suggests that grammatical items like ELLIPSIS require substantial memory from the early stages of training, and thus, our proposed method may not be optimal for learning such structures. Alternative workarounds, such as dynamically adjusting memory allocation or hybrid approaches, may be necessary to address this limitation.
@@ 5.2 Development of Feature Extraction Capabilities
Figure 3 visualizes the clustering structure of final-layer embeddings using t-SNE van der Maaten and Hinton (2008) for FILLER.GAP, a grammatical items where gradual memory expansion yielded significant performance improvements in both AO-CHILDES and Wikipedia datasets, as highlighted in the previous results (§4.3 and §5.1). In NoLimit (Figure 3a), the embedding clusters initially expand between Epoch 1 and Epoch 5, but by Epoch 10, they appear to contract and overlap more, suggesting a stagnation in representation learning. The clusters become less distinguishable, which may indicate a loss of diversity in the learned representations. In contrast, DynamicLimit-Exp (Figure 3b) maintains a more structured and progressive evolution of embeddings. The clusters remain well-separated throughout training, with clear distinctions between different epochs. This suggests that the model continuously refines its representations without excessive compression, preserving the diversity necessary for robust generalization.
To quantitatively analyze these differences, Table 4 reports key statistical measures, including entropy (distribution diversity) and mean Euclidean distance (inter-cluster separation). [^8] The appendix C shows how to calculate each measure. Regarding entropy, NoLimit shows a decreasing trend, reflecting reduced distribution diversity and potential over-clustering. In contrast, DynamicLimit-Exp preserves consistently higher entropy, indicating a balanced representation that avoids excessive compression. For mean Euclidean distance, NoLimit undergoes substantial change between Epoch 1 and Epoch 5 but stagnates thereafter, suggesting limited refinement. DynamicLimit-Exp, however, maintains large distances across epochs, reflecting continuous structural reorganization.
The differences also highlight the role of isotropy. NoLimit exhibits increasing anisotropy, with embedding clusters becoming overly compact by Epoch 10, which may hinder generalization. In contrast, DynamicLimit-Exp maintains a more isotropic distribution, as indicated by stable entropy, allowing for more flexible and structured representation learning. These findings align with recent work on syntactic smoothing, which suggests that reducing anisotropy enhances the ability to generalize across linguistic contexts Diehl Martinez et al. (2024). Thus, the increased isotropy observed in DynamicLimit-Exp provides strong evidence that gradual memory expansion facilitates structured representation learning and syntactic generalization. [^9] We also analyzed CASE, which exhibited the same trend as FILLER.GAP (as shown in Appendix D).
@@ 5.3 Influence of Input Stimulus Length
We analyze how sentence length affects the performance of NoLimit and DynamicLimit-Exp. To assess their adaptability, we created four Wikipedia-based datasets, each with 500,000 sentences in length ranges: [5,10], [11,50], [51,100], and [101,150].
The results in Table 5 reveal notable differences in model performance. For shorter sentences in the [5,10] range, NoLimit achieves slightly higher accuracy compared to DynamicLimit-Exp. However, in the [11,50] range, DynamicLimit-Exp significantly outperforms NoLimit, achieving 58.7 compared to 47.0. This suggests that DynamicLimit-Exp excels at handling moderately long sentences, likely due to its ability to dynamically adjust working memory. For longer sentences in the [51,100] and [101,150] ranges, DynamicLimit-Exp consistently outperforms NoLimit.
These findings highlight the benefits of dynamic working memory expansion in facilitating rule generalization and contextual adaptation across diverse sentence lengths. While NoLimit exhibits competitive performance on short sentences, its stagnation on longer sentences underscores its limited ability to generalize complex patterns. Conversely, DynamicLimit-Exp’s consistent performance across varying sentence lengths supports its suitability for grammatical items requiring the processing of both short and long contexts.
@@ 6 Conclusion
This study proposed a method for integrating the developmental trajectory of human working memory into the training process of LMs, inspired by the Less-is-More hypothesis. The proposed method, DynamicLimit-Exp, initially restricts working memory and gradually relaxes it exponentially during training. Experiments on both AO-CHILDES and Wikipedia showed that DynamicLimit-Exp improves grammatical learning efficiency compared to conventional methods without memory constraints or with static memory constraints. These findings suggest not only provide new approaches for developing data-efficient LMs but also offer indirect evidence supporting the CPH in human language acquisition.
@@ Acknowledgments
One limitation of this study is the constrained scale of the experimental setup. The primary goal of this study is to computationally replicate the CP in $\mathrm{L}_{1}$ acquisition, as discussed in cognitive science Lenneberg (1967); Fromkin et al. (1974); Curtiss (1977); Johnson and Newport (1989). Following previous studies Huebner et al. (2021); Haga et al. (2024), we designed the experiment to be as ecologically valid as possible by training an SLM using CDS. While this controlled setting allows for a more precise analysis and simulation of the Less-is-More hypothesis, it remains unclear how our findings contribute to the data efficiency of LLMs. The experimental results with Wikipedia (Table 2, 3, 5) provide a promising outlook in this direction, but further investigation with larger models and datasets is necessary to determine the effectiveness and limitations of the proposed approach.
@@
In this experiment, we investigated the replication of the CP effect in $\mathrm{L}_{1}$ acquisition using English. However, since the CP effect is observed across various languages Patkowski (1980); Johnson and Newport (1989), it remains to be tested whether the proposed approach is effective in multilingual environments. To our knowledge, there is currently no targeted syntactic evaluation specifically designed for CDS across different languages, such as Zorro. Zorro was developed based on BLiMP Warstadt et al. (2020), an adult-oriented targeted syntactic evaluation for English, and recent studies have proposed multilingual versions of BLiMP (e.g., JBLiMP Someya and Oseki (2023) for Japanese and CLiMP Xiang et al. (2021) for Chinese). Therefore, developing CDS-specific versions based on these multilingual BLiMPs could help address this limitation.
@@ Appendix A Details of the Training Configuration for the Base Models
Table 6 shows the training settings of the base model. For the experiment, a single NVIDIA RTX A5000 (24GB) GPU was used, and the training time for each run was approximately one hour.
@@
Maximum Sequence Length 32
@@
Optimizer AdamW
Optimizer Parameters $\beta=(0.9,0.999)$, $\epsilon=1e−08$
Tokenizer Trained on CHILDES
@@
Entropy Mean Distance
Epoch 1 5 10 1-5 5-10 1-10
NoLimit 5.30 5.23 5.30 75.47 12.26 87.62
DynamicLimit-Exp 5.29 5.30 5.34 59.91 37.68 97.59
Table 7: Embedded space analysis of NoLimit and DynamicLimit-Exp at each stage: cluster expansion, distribution diversity, and distribution distance.
@@ Appendix B Details of Grammatical Items in Zorro
Table 8: Explanation of each grammatical category in Zorro.
Table 8 shows the full list of grammatical categories in Zorro. Examples are taken from Table 5 in the original paper Huebner et al. (2021).
Appendix C Analysis of Distributional Changes in t-SNE Space Across Training Epochs
This section explains in detail the analysis of the entropy and average distance of embeddings projected into the t-SNE space for different learning epochs.
C.1 Entropy Calculation
To quantify the distribution of embeddings, a 2D histogram is constructed using a fixed grid ($50\times 50$ bins). The probability distribution $P$ is obtained by normalizing the histogram. The entropy is then computed as:
$H(P)=-\sum_{i}P_{i}\log P_{i},$ (4)
where $P_{i}$ is the probability of each bin. Higher entropy suggests a more uniform distribution, whereas lower entropy indicates clustering.
C.2 Mean Distance Between Epochs
To analyze shifts in embedding distributions across epochs, we compute the Euclidean distance between the mean embedding vectors of different epochs:
$D(X,Y)=\|\mu_{X}-\mu_{Y}\|,$ (5)
where $\mu_{X}$ and $\mu_{Y}$ are the mean vectors at different epochs. Larger distances imply greater shifts in the learned representation.
Appendix D Development of Feature Extraction Capabilities in CASE
(a) NoLimit
(b) DynamicLimit-Exp
Figure 4 visualizes the clustering structure of final layer embeddings using t-SNE for CASE. The embedding space visualizations reveal distinct patterns between NoLimit and DynamicLimit-Exp across training epochs. In NoLimit, the embedding clusters expand between Epoch 1 and Epoch 5 but contract significantly by Epoch 10, suggesting stagnation in representation learning. In contrast, DynamicLimit-Exp maintains structured evolution throughout training, with well-separated clusters that reflect progressive refinement.